OpenStudy (anonymous):

tan θ sin θ + sin θ = 0 Solve the given equation.

OpenStudy (anonymous):

First step is to factor the left side.

OpenStudy (anonymous):

Can you do that?

OpenStudy (anonymous):

I'm not really sure how. This is the first type of this particular problem on my homework assignment.

OpenStudy (anonymous):

How would you factor ax+x?

OpenStudy (anonymous):

Factor out the X

OpenStudy (anonymous):

Sin Theta(Tan Theta + 1) ?

OpenStudy (anonymous):

right

OpenStudy (anonymous):

Do you know what's next?

OpenStudy (anonymous):

Try to get it as a single trig identity? I'm not sure...

OpenStudy (anonymous):

you have a product of two things: sin(theta) and tan(theta + 1) and that product equals 0, so at least one factor is 0.

OpenStudy (anonymous):

So sin(theta)=0 or tan(theta)+1=0 Similar to how you would solve (x+1)(x+2)=0 -> x+1=0 or x+2=0 for example.

OpenStudy (anonymous):

So I just find where they equal 0?

OpenStudy (anonymous):

Sin theta = 0 at 2pi and pi, and tan theta = -1 at 7pi/4 ad 5pi/4

OpenStudy (anonymous):

And then I just add on 2pi K and pi K for where Sin theta = 0 and where Tan theta = -1?

OpenStudy (anonymous):

I think..

OpenStudy (anonymous):

you're right about the sine part, but tan(5pi/4)=1

OpenStudy (anonymous):

oooo 3pi/4, got the circle mixed up

OpenStudy (anonymous):

So would the solution be 2pi+2pi k, pi + 2pi K, 7pi/4 + pi k, and 3pi/4 +pi k?

OpenStudy (anonymous):

yes, you can simplify that though.

OpenStudy (anonymous):

Alright I got the question right

OpenStudy (anonymous):

Thanks a lot!

OpenStudy (anonymous):

I feel compelled to write down the simplified answer: x=k pi or x= k pi -1/4 pi

OpenStudy (anonymous):

Gotcha, My professor allows us to leave our answers unsimplified. Thanks again.

OpenStudy (anonymous):

ok then, you're welcome.